What is the common difference in the arithmetic sequence 3.7,4.4,4.9,5.5,6.1,...
Answer by Edwin McCravy(20056)

Apparently the second term is mistyped as 4.4.  For that to be 
an arithmetic sequence the second term would have to be 4.3.

3.7, 4.3, 4.9, 5.5, 6.1, ...

Then 

2nd term - 1st term = 4.3-3.7 = 0.6
3rd term - 2nd term = 4.9-4.3 = 0.6
4th term - 3rd term = 5.5-4.9 = 0.6
5th term - 4th term = 6.1-5.5 = 0.6

So the common difference is 0.6

Edwin